OpenStudy (anonymous):

-2/5x=-4 solve for the given variable

jimthompson5910 (jim_thompson5910):

Start by multiplying both sides by 5 to go from -2/5x=-4 to 5*(-2/5)x=-4*5 -2x = -20

jimthompson5910 (jim_thompson5910):

Then what's next?

OpenStudy (anonymous):

multiply both sides by the reciprocal of -2/5 x=-4(-5/2) x=-10

jimthompson5910 (jim_thompson5910):

answer should be positive yummydum

OpenStudy (anonymous):

oh right sorry..thanks @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

you're welcome